OpenStudy (anonymous):

Add and simplify 5/49z^3+1/35z^2=

OpenStudy (saifoo.khan):

First step, find the LCM, 245z^3 Then solve both fractions, \[\frac{7z+25}{245z^3}\]

hero (hero):

You don't need to find an LCD to solve it

OpenStudy (saifoo.khan):

Yes we do actually. b/c it's \[\frac{5}{49z^3} + \frac{1}{35z^2} \]

hero (hero):

She put an equal sign there. It threw me off.

hero (hero):

You're still double zero

OpenStudy (saifoo.khan):

That proves that you are blind and illiterate. That's 'o' not '0'.

hero (hero):

double zer'o'

OpenStudy (saifoo.khan):

Blind and illiterate.

hero (hero):

I don't see a 0 in Hero either. I guess you're just as blind and illiterate